NAOJ GW Elog Logbook 3.2

[Hugo, Marc]
Today we wanted to restart birefringence measurement with PCI and longer average.
We found out that the motorized rotator is broken (sounds but no motion).
We ordered 2 new ones.

In the spare OPO box in my office there is a spare retaining ring but for whatever reason no spare front plate PZT holder.
Since in the actual assembly the inner retaining ring only needs to be screwed down to the bottom of the wire extraction groove (see fig 1, 2, 3 of 2690), maybe I should just screw in the spare without replacing the front plate. I honestly have no idea why this PZT holder hole is even cut that deep.

Current ATC IR beam profile after the 100mm/100mm telescope.

[Marc, Shalika]
There was an electrical shutdown in Mitaka campus on Saturday due to the heavy storm.
We recovered the following :
LC table : pc, laser and electronics
PCI table : pc, laser and translation stage and electronics
However, I misoperate the translation stage that hit the polarizer mount installed for birefringence measurement.
The good recovery procedure is the following:
- remove all components on the translation stage
- using the manual controller on the side of the imaging unit box, moved the translation stage in X to be above any components in the injection bench.
- using zaber, home all motors
- use manual controller to move in reasonable Z before moving in a reasonable X. here reasonable meaning far enough from both injection and readout benches.
- close zaber and restart LabView
- FDS : recovered the reference 'gps' signal, recovered the temperature controls and all pcs. Because of fire safety inspection, we did not check fds laser and we also need to recover the homodyne power supply, AOM controller and restart vacuum pumps
- ATC cleanroom : restarted the fans
- cryogenic cavity : pumps seems stopped but it seems somebody else will check
[Hugo, Marc]
Today we wanted to restart birefringence measurement with PCI and longer average.
We found out that the motorized rotator is broken (sounds but no motion).
We ordered 2 new ones.

The nonlinear crystal PPKTP is set up on a copper plane with a TC cooler below the crystal, these components are on a one-dimensional motion stage.
The TC cooler is monitored via a temperature sensor (TT2-10KC8-3) and controlled by a temperature controller (THORLABS TC300).
The linear motion stage can be driven directly through an electronic controller (PI E-873 Q-Motion Controller) and software installed on Windows 10 or higher.

"Birefringence Map with Camera.vi" modified to take into account std and 100 averaging
file will save in format
X, Y, Azimuth, std azi, Ellipticity, std ell, Power, s1, s2, s3

The bowtie VOPO was completely assembled. It looks quite nice, and not so difficult to work with. The crystal and mirror positions are well defined in the assembly. I haven't seen the green alignment in action, but it seems much more convenient to be able to actually see the beam path inside the OPO. Considering also the advantages in controllability, the lack of need for a GRMC ($) and ppol laser ($$$), and our regular collaboration with NTHU, I aim to use this type of OPO in all future quantum optics experiments at NAOJ or elsewhere. I asked Chien Ming and Hsun Chung to give me lots of documentation. I also gave Chien Ming the new elog link to give more details, since I couldn't be there the entire time.
I briefly came back to reality with the TAMA linear OPO. The PZT retaining ring is completely stuck. Even using gorilla force with Thorlabs SPW801 wrench I cannot move it at all. It seems more likely that I would just rip out the metal (fig 1). I tried to check the OPO but it's a bit hard to focus (fig 2). I suppose I will have to give up and use the spare front plate (in a box in my office) and spare retaining ring (somewhere...). I will need to clean them with the ultrasonic bath in the ATC.
With permission of Aso-san I took the small First Contact bottle from the speedmeter experiment, though he mentioned that he didn't know where Nishino-kun got it from originally. I spent a lot of time looking for the tabs used to peel it. Eventually I found a pack of them hiding at the back of the ATC cleanroom. I somewhat skilllessly applied a layer to the TAMA OPO incoupler convex face (fig 3). From experience last time, this won't remove the O-ring residue, but I just want to get the dust off the centre. Next time I will do the concave face.
In the spare OPO box in my office there is a spare retaining ring but for whatever reason no spare front plate PZT holder.
Since in the actual assembly the inner retaining ring only needs to be screwed down to the bottom of the wire extraction groove (see fig 1, 2, 3 of 2690), maybe I should just screw in the spare without replacing the front plate. I honestly have no idea why this PZT holder hole is even cut that deep.

I took the following:
Mirror mounts - one from scattering table, two from FC cleanroom
IR steering mirrors - one pack of 9 from the KFC spare parts box
Fiber PD power supply - To TAMA
It seems Newport's SW-OM adjustable spanner wrench may be more suited to getting rid of the TAMA OPO retaining ring. There wasn't any appropriate tool in the TAMA OPO spare parts box. Thorlabs sells SPW-908 that looks like the exact kind of thing to unscrew the retaining ring, but conveniently the unscrewing things are 1mm too far apart.

I will set up blackout curtain near the cryostat optical table to section off the back side of ATC lab (figure-1). The plan is to move the long Al bar closer to the AC unit and hang the curtain on the entire length of the rod (figure-2). I measured various dimensions for this work.
During this work, I noticed the items on the optical table were moved. If you move stuff on the table, an email or verbal confrimation would be appericated.
Hirata gave me some drill bits yesterday. Unfortuantely. these drill bits are too small. If I can't find one in ATC I will order them online. If some has HEX5 bits for Makita drill, similar to the bit in the attached picture please let me know.

Hsun-Chung, Chien-Ming, Michael
We took from TAMA the Qubig EOM box as well as a Sigma-type 30mm polarizer mount and a regular steering mirror mount. We also brought the VOPO equipment to the ATC cleanroom.
I also brought from FC cleanroom an adjustable-width key to try and unscrew the PZT spacer from the assembly (fig 6 of 3651) but the necessary force is ridiculously tight and I ended up just scratching the brass face again. I need to look for a better tool. Maybe this is why all serious squeezing experiments use the Advanced LIGO bowtie design instead.

Logan, Michael
We inspected the 0.5in incoupler mirror of the OPO cavity using a strong flashlight. I took very close up photos with my macro lens (fig 1). There is a lot of o-ring residue which is unavoidable. Also a bit of dust. Maybe won't hurt to put First Contact.
We took out the OPO PZT and inspected (fig 2, 3). The new one looks a bit cleaner (fig 4, 5). Maybe the electrode on the old one is burnt out? I don't really know but maybe should carefully confirm the polarity of the PZT before soldering.
When replacing the PZT the retaining ring was screwed in too far and was hard to unscrew with stuff in the ATC room (fig 6). I'm pretty sure there is a proper tool in TAMA. That said I also think this OPO spacer design is very bad.


50L, 5/6 WITHOUT SCREWS STRAIN
average = 15 ; waits time =10 000ms
air r = 3 ; sample r = 10
center x= 323.565; y= 154.375 ; z=60
azy=75
air= Wed, Jul 10, 2024 9-10-19 AM.txt
sample: Wed, Jul 10, 2024 9-18-28 AM.txt
azy=60
air=Wed, Jul 10, 2024 10-32-55 AM.txt
sample=Wed, Jul 10, 2024 10-42-45 AM.txt
azy=45
air = Wed, Jul 10, 2024 11-58-29 AM.txt
sample=Wed, Jul 10, 2024 12-36-48 PM.txt
azy=30
air=Wed, Jul 10, 2024 1-51-28 PM.txt
sample=Wed, Jul 10, 2024 1-58-31 PM.txt
azy=15
air=Wed, Jul 10, 2024 3-43-55 PM.txt
sample=Wed, Jul 10, 2024 3-52-17 PM.txt
azy=0
air = Wed, Jul 10, 2024 5-06-51 PM.txt
sample=Wed, Jul 10, 2024 5-25-48 PM.txt

After creating different maps with averages of 15, 50, and 100, we are still encountering issues with the homogeneity of the Jones matrix. We suspect there may be problems with the LabVIEW code conducting measurements during the translation stage movement. Therefore, we are trying a new measurement approach by increasing the wait time.
50L, 5/6 WITHOUT SCREWS STRAIN
average = 15 ; waits time =1000ms
air r = 3 ; sample r = 3
center x= 323.565; y= 154.375 ; z=60
azy= 0
air : Tue, Jul 9, 2024 6-37-07 PM.txt
sample : Tue, Jul 9, 2024 6-40-33 PM.txt
azy=15
air: Tue, Jul 9, 2024 6-43-29 PM.txt
sample : Tue, Jul 9, 2024 6-46-46 PM.txt
azy=30
air: Tue, Jul 9, 2024 6-49-47 PM.txt
sample: Tue, Jul 9, 2024 6-52-26 PM.txt
azy=45
air :Tue, Jul 9, 2024 6-55-41 PM.txt
sample :Tue, Jul 9, 2024 6-58-18 PM.txt
azy=60
air:Tue, Jul 9, 2024 7-01-33 PM.txt
sample:Tue, Jul 9, 2024 7-04-16 PM.txt
azy=75
air=Tue, Jul 9, 2024 7-08-22 PM.txt
sample =Tue, Jul 9, 2024 7-11-16 PM.txt

50L, 5/6 WITHOUT SCREWS STRAIN
average = 100
air r = 3 ; sample r = 13
center x= 323.565; y= 154.375 ; z=60
azy=75
air=Mon, Jul 8, 2024 7-33-54 AM.txt
sample = Mon, Jul 8, 2024 7-44-50 AM.txt
azy=60:
air = Mon, Jul 8, 2024 10-52-11 AM.txt
sample = Mon, Jul 8, 2024 12-48-55 PM.txt
azy=45
air=Mon, Jul 8, 2024 3-55-57 PM.txt
sample : Mon, Jul 8, 2024 4-06-11 PM.txt
azy=30
air= Mon, Jul 8, 2024 8-01-26 PM.txt
sample : Mon, Jul 8, 2024 8-12-17 PM.txt
azy=15
air = Tue, Jul 9, 2024 7-10-43 AM.txt
sample :Tue, Jul 9, 2024 10-29-25 AM.txt
azy = 0
air = Tue, Jul 9, 2024 1-37-48 PM.txt
sample = Tue, Jul 9, 2024 1-48-34 PM.txt

I took from general TAMA storage a couple of polarizer mounts (one Thorlabs an done Newport), as well as rail mounts and ND filters from the squeezing clean room.
Last Friday I checked a bit the beam size based roughly on a recalculated version of the OPO MMT from Logan. While in principle we could design the beam to be a little bit too small (~20um) and then fine tune it upwards in size, in reality I was having trouble approaching the required 25 um radius. Somehow I seem to be having more trouble with this than I expected. Last time it "just worked" straight out of JamMT. Sometimes I see a very small beam on the sensor card but the detector tells me it is 0.5mm diameter as well.
I suspected a bit that we might be saturating the detector but according to the specs that doesn't seem to be the case (approximate intensity = 0.002 W /(pi*(0.020 mm)^2 = 1.6 W/mm^2 - max dynamic range 10W/mm^2, damage threshold 200 W/mm^2). Still, I brought stronger ND filtering to check anyway.
I disassembled the OPO input optics assembly and took out the input mirror. The PZT retaining rings seems slightly dusty. I tried to check with my weak phone light if the input mirror was dirty but I couldn't see anything. Like last time, there seems to be some residue from the O-ring where it is pressing against the input coupler clamp. I left the input mirror on top of some Alpha wipes with Akutsu-san's stuff but then saw that they were a bit dirty, there is even a mosquito on the cloth. Also I have some soldering gear (iron, heater) sitting outside the ATC cleanroom but I won't be available to finish the soldering job until Thursday unfortunately.

We get some issue with the analysis maybe due to the low averaging so I do the same with an more importante averaging.
50L, 5/6 WITHOUT SCREWS STRAIN
average = 50
air r = 3 ; sample r = 13
center x= 323.565; y= 154.375 ; z=60
azy= 0
air : Fri, Jul 5, 2024 12-06-42 PM.txt
sample : Fri, Jul 5, 2024 12-37-36 PM.txt
azy = 15
air : Fri, Jul 5, 2024 2-17-08 PM.txt
sample : Fri, Jul 5, 2024 2-22-40 PM.txt
azy=30
air = Sun, Jul 7, 2024 8-15-29 AM.txt
sample = Sun, Jul 7, 2024 8-21-47 AM.txt
azy=45
air=Sun, Jul 7, 2024 4-04-18 PM.txt
sample =Sun, Jul 7, 2024 4-10-02 PM.txt
azy = 60
air = Sun, Jul 7, 2024 6-09-54 PM.txt
sample = Sun, Jul 7, 2024 6-16-48 PM.txt
azy=75
air = Sun, Jul 7, 2024 8-01-19 PM.txt
sample : Sun, Jul 7, 2024 8-07-07 PM.txt

@ 169 deg
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Coating measurement\AlGaAs\20240701\169 deg\Fri, Jul 5, 2024 3-25-49 PM.txt

@ 167 deg
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Coating measurement\AlGaAs\20240701\167 deg\Fri, Jul 5, 2024 1-23-23 PM.txt

error last raw replace by : "sample : Fri, Jul 5, 2024 9-10-47 AM.txt"